The sum of 2 consecutive even numbers is greater than or equal to 158. What are the
minimum values of the two numbers?

Answer by ikleyn(52798)

.

    n + (n+2) >= 158


    2n        >= 158 - 2 = 156


     n        >= 156/2 = 78.


The minimum value of the two numbers is 78.    ANSWER

Solved.